The sum of five numbers in arithmetic progression is 25 and of their square is 165. Find the numbers.
1
Expert's answer
2020-06-22T18:16:07-0400
"a_1+a_1+d+a_1+2d+a_1+3d+a_1+4d=25"

"a_1^2+(a_1+d)^2+(a_1+2d)^2+""+(a_1+3d)^2+(a_1+4d)^2=165"

"a_1=5-2d"

"(5-2d)^2+(5-d)^2+(5)^2+(5+d)^2+(5+2d)^2=165"

"25-20d+4d^2+25-10d+d^2+25+""+25+10d+d^2+25+20d+4d^2=165"

"10d^2=40"

"d^2=4"

"d_1=-2, d_2=2"

"d=-2"

"a_1=5-2(-2)=9"

The numbers are "9,7,5,3,1."


"d=2"

"a_1=5-2(2)=1"

The numbers are "1,3,5,7,9."